# Frozen Transformers in Language Models Are Effective Visual Encoder Layers

### v1

@@ Abstract
This paper reveals that large language models (LLMs), despite being trained solely on textual data, are surprisingly strong encoders for purely visual tasks in the absence of language. Even more intriguingly, this can be achieved by a simple yet previously overlooked strategy – employing a frozen transformer block from pre-trained LLMs as a constituent encoder layer to directly process visual tokens. Our work pushes the boundaries of leveraging LLMs for computer vision tasks, significantly departing from conventional practices that typically necessitate a multi-modal vision-language setup with associated language prompts, inputs, or outputs. We demonstrate that our approach consistently enhances performance across a diverse range of tasks, encompassing pure 2D and 3D visual recognition tasks (e.g., image and point cloud classification), temporal modeling tasks (e.g., action recognition), non-semantic tasks (e.g., motion forecasting), and multi-modal tasks (e.g., 2D/3D visual question answering and image-text retrieval). Such improvements are a general phenomenon, applicable to various types of LLMs (e.g., LLaMA and OPT) and different LLM transformer blocks. We additionally propose the information filtering hypothesis to explain the effectiveness of pre-trained LLMs in visual encoding – the pre-trained LLM transformer blocks discern informative visual tokens and further amplify their effect. This hypothesis is empirically supported by the observation that the feature activation, after training with LLM transformer blocks, exhibits a stronger focus on relevant regions. We hope that our work inspires new perspectives on utilizing LLMs and deepening our understanding of their underlying mechanisms. Code is available at https://github.com/ziqipang/LM4VisualEncoding.
@@ 1 Introduction
Large language models (LLMs), trained on massive amounts of textual data, have recently demonstrated remarkable potential across various tasks, extending beyond their original linguistic domain. For example, in the field of computer vision, LLMs exhibit the ability to interact with visual tokens and decode them into tokenized output. This is commonly achieved in a multi-modal vision-language framework that incorporates the language modality, as exemplified by either projecting visual tokens to LLMs via linear layers (Koh et al., 2023; Lin et al., 2023; Merullo et al., 2023; Schwettmann et al., 2023) or employing cross-attention mechanisms between visual and language tokens (Alayrac et al., 2022; Li et al., 2022; 2023; Wang et al., 2023). As we explore the limits of utilizing LLMs for computer vision tasks, an interesting question arises: can LLMs effectively handle tasks that are exclusively visual, without any reliance on language?
@@
However, one crucial question remains: why are LLMs effective in visual encoding, given that they have been exclusively trained on texts and have never encountered visual data? To this end, we propose the information filtering hypothesis: the pre-trained LLM transformer blocks discern informative visual tokens and further amplify their contribution to the latent representation. This hypothesis stems from our observation across multiple tasks, where the feature activation consistently exhibits a stronger focus on relevant regions, after integrating the frozen LLM transformer blocks.
@@
We hope that our work will drawn attention to the intriguing application of employing LLM transformers as versatile encoders, not only for visual inputs but potentially also for other modalities. Additionally, we hope to inspire new perspectives on understanding LLMs/VLMs.
@@ 2 Related Work
Large language models.   Pretraining transformers (Vaswani et al., 2017) with masked token prediction facilitates the generalizability of language models (LMs) to various tasks, represented by BERT (Kenton & Toutanova, 2019). Later on, larger models at scale are proposed guided by the scaling law (Kaplan et al., 2020), such as GPT (Brown et al., 2020), LLaMA (Touvron et al., 2023), OPT (Zhang et al., 2022), BLOOM (Scao et al., 2022), GLM (Zeng et al., 2022), PaLM (Chowdhery et al., 2022) and models enhanced with additional tuning steps (Taori et al., 2023; Zhang et al., 2023a), etc. These large models with tens of billions of parameters unlock the intriguing ability of in-context learning and excellent zero-shot performance on various tasks. Our work highlights the interesting discovery that the transformer blocks in such large language models (LLMs) are able to interact with visual data and enhance a wide spectrum of computer vision tasks.
@@
Language models for visual tasks.   LMs are mostly used as language encoders for vision-language models (VLMs) (Dou et al., 2022; Kim et al., 2021) or image-text pretraining (Radford et al., 2021) before the emergence of LLMs. Such applications are still widely existing in scenarios like text-to-image generation (Rombach et al., 2022) and open-world perception (Kirillov et al., 2023). After the creation of LLMs, their code generation ability encourages flexibly combining the vision algorithms for user queries, (Shen et al., 2023; Surís et al., 2023), represented by visual programming (Gupta & Kembhavi, 2023). In addition, the strong ability of LLMs has also elicited using them as generalizable decoders, i.e., translating the latent feature into output tokens (Wang et al., 2023; Zhang et al., 2023b). These frameworks commonly project visual features to the input layer of LLMs directly (Guo et al., 2023; Koh et al., 2023; Lin et al., 2023; Merullo et al., 2023) or use the structures of latent bottleneck (Jaegle et al., 2021) to further encode visual tokens (Alayrac et al., 2022; Hong et al., 2023; Li et al., 2022; 2023; Wang et al., 2023). Our exploration reveals the potential of considering the transformer blocks in LLMs as general-purpose encoders for visual data, as opposed to the previous usages of either pure encoders for text embeddings or decoders for tokenized outputs.
@@
Interpreting neural networks.   Understanding neural networks begins by visualizing the convolutional patterns in low-level layers (Erhan et al., 2009). With a deeper interest in semantics, attribution-based methods like Grad-CAM (Selvaraju et al., 2017) further analyze the contribution of neurons for a certain class. The network dissection studies (Bau et al., 2017; Zhou et al., 2018) also discover that neural network units correspond to semantic concepts. In addition to post-hoc interpretation, Böhle et al. (2022; 2023); Yu et al. (2023) build “white-box” models for clear interpretability. For LLMs, researchers discover that the knowledge is mainly located at the MLPs in feedforward networks (FFN) (Dai et al., 2022; Geva et al., 2020; Meng et al., 2022a; b), and sometimes correspond to visual concepts (Schwettmann et al., 2023). Compared to them, we study a new scenario of why a pre-trained LLM transformer can benefit visual encoding and propose the information filtering hypothesis.
3 Method: Frozen LLM Transformer for Visual Encoding
@@
Framework design.   We formally introduce using a pre-trained LLM transformer as a visual encoder layer shown in Fig. 0(a). Without loss of generality, we consider a neural network that maps an input $x$ to latent representation $z$ and predicts labels $y$ with an encoder $\mathbf{F}_{E}$ and decoder $\mathbf{F}_{D}$,
@@
In the training stage, the pretrained transformer $\mathbf{F}_{LM}$ remains frozen, as in the pseudo-code of Fig. 0(b), while all the other modules are trained normally, including $\mathbf{F}^{1}_{L}$ and $\mathbf{F}^{2}_{L}$.
Comparison with vision-language models.   Our approach appears similar to recent vision-language models (VLMs) at the first glance, such as  Lin et al. (2023), FROMAGe (Koh et al., 2023), and LiMBeR (Merullo et al., 2023), where linear layers directly project visual features to the input space of LLMs. However, our approach is different, because the linear layer $\mathbf{F}_{L}^{1}$ does not necessarily align the visual representation $z$ into the language space. Concretely, this is reflected in three aspects: (1) Independence of visual pre-training. Our paradigm supports training-from-scratch without relying on pre-trained visual encoders like CLIP (Radford et al., 2021). (2) Independence of language. Our framework can function without language-based input or prompts, and it is applicable for general visual representation learning instead of only vision-language tasks. (3) Independence of transformer blocks. Previous VLMs treat an LLM as a coherent module, while our framework separates each transformer block as an independent layer for visual encoding.
Comparison with LLMs.   We substantially change the behaviors of LLM transformers, due to the distinct formats between visual and textual data. (1) Attention mask. LLMs commonly utilize auto-regressive masks to mimic the order of text generation. However, the tokens in visual data come all at once, such as the image tokens of the cat (Fig. 0(a)). So we abandon auto-regressive attention masks and only use attention masks to indicate the padded tokens. (2) Positional embedding. The positional embedding in LLMs, e.g., rotary positional embedding (Su et al., 2021) in LLaMA, is not a common option for visual encoders. Therefore, we remove the positional embeddings of LLMs for simplicity and consistency with the original visual backbones. Considering the importance of attention masks and positional embeddings in transformers, it is surprising in hindsight that our framework has a positive influence on visual tasks.
@@ 4 Applicability of LLM Transformers for Visual Tasks
We instantiate our framework to various tasks and observe the wide applicability of pretrained LLM transformers. Our experiments cover 2D (image classification) and 3D (point cloud classification), single-frame and multi-frame (action recognition), semantics and motion (motion forecasting), and involving linguistic input or not (2D and 3D vision-language tasks). By default, we adopt the last transformer block from LLaMA-7B (Touvron et al., 2023). Our framework achieves consistent and significant improvements across these tasks following the standards of prior work. More analysis on our designs is in Sec. 5.
@@ 4.1 Image Classification
Task statement.   Image classification is the most common challenge for representation learning. We conduct experiments on ImageNet1k (Deng et al., 2009) for accuracy, and additionally evaluate on robustness benchmarks: corrupted images from ImageNet-C (Hendrycks & Dietterich, 2018), natural adversarial images from ImageNet-A (Hendrycks et al., 2021b), and out-of-distribution images from ImageNet-SK (Wang et al., 2019) and ImageNet-R (Hendrycks et al., 2021a).
Design.   Without loss of generality, we select ViT (Dosovitskiy et al., 2021) due to its wide use and native support for transformers. Following the notation in Eqn. 2, the encoder $\mathbf{F}_{E}$ is the set of self-attention transformer blocks in ViT and the decoder $\mathbf{F}_{D}$ is the linear classifier, respectively. The LLaMA transformer is inserted between the encoder and decoder according to the illustration in Fig. 0(a). We train both the baseline ViT models and ViT-LLaMA from scratch following the same configuration of DeiT (Touvron et al., 2021). Please refer to Sec. C.1 for implementation details.
Analysis.   The accuracy of ViT models consistently improves after incorporating the frozen LLaMA transformer block as in Table 1, including both the accuracy on clean ImageNet images and robustness on corrupted, adversarial, and out-of-distribution images. Our further experiments validate that the improvement is closely related to the LLM transformer instead of the sole consequence of an increased model capacity. Please refer to Sec. 5.1 for implementation details. We further explain the source of improvement with our information filtering hypothesis in Sec. 6.
4.2 Point Cloud Recognition
Task statement.   Point cloud classification handles a fundamentally different modality compared to images. The models predict the labels by processing unordered 3D points and understanding the local-global geometry. Our experiments cover two common datasets: ScanObjectNN (Uy et al., 2019) and ModelNet40 (Goyal et al., 2021). ScanObjectNN contains three splits: background points (BG), foreground (OBJ), and clipped (T50) points. For ModelNet40, we experiment with different densities (1k, 4k, 8k) of points.
@@ 4.2 Point Cloud Recognition
Table 2: LLM transformer improves point cloud recognition.
Design.   We adopt Point-BERT (Yu et al., 2021) as our base framework. Since Point-BERT first pretrains on ShapeNet (Chang et al., 2015) then finetunes on downstream point cloud classification, we directly load the pretrained parameters on ShapeNet in our experiments. Then we append the LLaMA transformer after its final attention block before finetuning on point cloud classification datasets. The details are in Sec. C.2.
Analysis.   As shown in Table 2, our approach improves the accuracy of point cloud classification and further supports the applicability of using a frozen LLM transformer as a visual encoding layer. Note that the accuracy slightly drops on ModelNet40 with 1k and 4k points, due to the saturation and $\sim$0.2variance of performance on ModelNet40, which is also analyzed in Ma et al. (2022). However, with an increased number of points (8k), the improvement of the LLaMA transformer is noticeable on ModelNet40. More importantly, on the more challenging ScanObjectNN, our approach improves the accuracy consistently and significantly, especially the most difficult splits of background (BG) and clipped (T50) point clouds. This experiment also shows that our framework is compatible with finetuning setups, in addition to training-from-scratch scenarios in Sec. 4.1.
@@ 4.3 Action Recognition
Task statement.   For the video modality, we apply the pretrained LLM transformer block to action recognition, where the algorithm predicts the action labels by processing video clips lasting a few seconds. We choose the benchmark of “Something-something-v2” (SSv2) (Goyal et al., 2017b) for evaluation, because it highlights the challenge of understanding cross-frame movement, instead of relying on single-frame semantics.
@@
Table 3: Frozen LLM transformer improves action recognition.
Design.   We follow VideoMAE (Tong et al., 2022) and adopt the simple yet effective ViT backbones for action recognition. Different from patches of tokens in 2D images, the video tokens are cubes spanning both spatially and temporally. When adding the LLaMA transformer, we place it behind the last self-attention block in ViT, identical to Fig. 0(a). According to the practice in VideoMAE, our model training involves two steps: (1) initialize the parameters of ViT transformers from VideoMAE’s MAE (He et al., 2022) pre-training; (2) add the LLM transformer behind the ViT blocks and then finetune on the SSv2 dataset. More implementation details are in Sec. C.3.
Analysis.   The LLaMA transformer enhances the accuracy for both ViT-S and ViT-B in Table 3, supporting the applicability of our framework for spatial-temporal information in videos. To clarify, our baseline accuracy is lower than VideoMAE’s original paper, because VideoMAE used 32/64 GPUs to enable a larger batch size than our computational resources. Nonetheless, we control the settings identical between ViT and ViT-LLaMA for a fair comparison and indicate the positive effects of using LLM transformers.

### v2

@@ Abstract
This paper reveals that large language models (LLMs), despite being trained solely on text data, are surprisingly strong encoders for purely visual tasks in the absence of language. Even more intriguingly, this can be achieved by a simple yet previously overlooked strategy – employing a frozen transformer block from pre-trained LLMs as a constituent encoder layer to directly process visual tokens. Our work pushes the boundaries of leveraging LLMs for computer vision tasks, significantly departing from conventional practices that typically necessitate a multi-modal vision-language setup with associated language prompts, inputs, or outputs. We demonstrate that our approach consistently enhances performance across a diverse range of tasks, encompassing purely 2D and 3D visual recognition tasks (e.g., image and point cloud classification), temporal modeling tasks (e.g., action recognition), non-semantic tasks (e.g., motion forecasting), and multi-modal tasks (e.g., 2D/3D visual question answering and image-text retrieval). Such improvements are a general phenomenon, applicable to various types of LLMs (e.g., LLaMA and OPT) and different LLM transformer blocks.
We additionally propose the information filtering hypothesis to explain the effectiveness of pre-trained LLMs in visual encoding – the pre-trained LLM transformer blocks discern informative visual tokens and further amplify their effect. This hypothesis is empirically supported by the observation that the feature activation, after training with LLM transformer blocks, exhibits a stronger focus on relevant regions. We hope that our work inspires new perspectives on utilizing LLMs and deepening our understanding of their underlying mechanisms.
@@ 1 Introduction
Large language models (LLMs), trained on massive amounts of text data, have recently demonstrated remarkable potential across various tasks, extending beyond their original linguistic domain. For example, in the field of computer vision, LLMs exhibit the ability to interact with visual tokens and decode them into tokenized output. This is commonly achieved in a multi-modal vision-language framework that incorporates the language modality, as exemplified by either projecting visual tokens to LLMs via linear layers (Koh et al., 2023; Lin et al., 2023; Merullo et al., 2023; Schwettmann et al., 2023) or employing cross-attention mechanisms between visual and language tokens (Alayrac et al., 2022; Li et al., 2022; 2023; Wang et al., 2023). As we explore the limits of utilizing LLMs for computer vision tasks, an interesting question arises: can LLMs effectively handle tasks that are exclusively visual, without any reliance on language?
@@
However, one crucial question remains: why are LLMs effective in visual encoding, given that they have been exclusively trained on text and have never encountered visual input? To this end, we propose the information filtering hypothesis: the pre-trained LLM transformer blocks discern informative visual tokens and further amplify their contribution to the latent representation. This hypothesis stems from our observation across multiple tasks, where the feature activation consistently exhibits a stronger focus on relevant regions, after integrating the frozen LLM transformer blocks.
@@
We hope that our work will drawn attention to the intriguing application of employing LLM transformers as versatile encoders, not only for visual inputs but potentially also for other modalities. Additionally, we hope to inspire new perspectives on understanding LLMs and VLMs.
@@ 2 Related Work
Large language models.   Pre-training transformers (Vaswani et al., 2017) with masked token prediction facilitates the generalizability of language models (LMs) to various tasks, represented by BERT (Kenton & Toutanova, 2019). Later on, larger models at scale are proposed guided by the scaling law (Kaplan et al., 2020), such as GPT (Brown et al., 2020), LLaMA (Touvron et al., 2023), OPT (Zhang et al., 2022), etc. These large models with tens of billions of parameters unlock the intriguing ability of in-context learning and excellent zero-shot performance on various tasks. Our work highlights the interesting discovery that the transformer blocks in such large language models (LLMs) are able to interact with visual data and enhance a wide spectrum of computer vision tasks.
@@
Language models for visual tasks.   LMs are mostly used as text encoders for vision-language models (VLMs) (Dou et al., 2022; Kim et al., 2021) or image-text pre-training (Radford et al., 2021) before the emergence of LLMs. After the creation of LLMs, their code generation ability encourages flexibly combining the vision algorithms for user queries, represented by visual programming (Gupta & Kembhavi, 2023). In addition, the strong ability of LLMs has also elicited using them as generalizable decoders, i.e., translating the latent feature into output tokens. These frameworks commonly project visual features to the input layer of LLMs directly (Guo et al., 2023; Koh et al., 2023; Lin et al., 2023; Merullo et al., 2023; Chen et al., 2023) or use the structures of the latent bottleneck (Jaegle et al., 2021) to further encode visual tokens (Alayrac et al., 2022; Hong et al., 2023; Li et al., 2022; 2023; Wang et al., 2023). Our exploration reveals the potential of considering the transformer blocks in LLMs as general-purpose encoders for visual data, as opposed to the previous usages of either pure encoders for text embeddings or decoders for tokenized outputs.
@@
Interpreting neural networks.   Understanding neural networks begins by visualizing the convolutional patterns in low-level layers (Erhan et al., 2009). With a deeper interest in semantics, attribution-based methods like Grad-CAM (Selvaraju et al., 2017) further analyze the contribution of neurons for a certain class. Network dissection  (Bau et al., 2017; Zhou et al., 2018) also discovers that neural network units correspond to semantic concepts. For LLMs, researchers find that the knowledge is mainly located at the linear layers in feedforward networks (FFN) (Dai et al., 2022; Geva et al., 2020; Meng et al., 2022a; b), and corresponds to visual concepts (Schwettmann et al., 2023). Compared with them, we study a new scenario of why a pre-trained LLM transformer can benefit visual encoding and propose the information filtering hypothesis.
3 Method: Frozen LLM Transformers for Visual Encoding
@@
Framework design.   We formally introduce using a pre-trained LLM transformer as a visual encoder layer shown in Fig. 1(a). Without loss of generality, we consider a neural network that maps input $x$ to latent representation $z$ and predicts labels $y$ with an encoder $\mathbf{F}_{E}$ and a decoder $\mathbf{F}_{D}$,
@@
In the training stage, the pre-trained transformer $\mathbf{F}_{LM}$ remains frozen, as in the pseudo-code of Fig. 1(b), while all the other modules are trained normally, including $\mathbf{F}^{1}_{L}$ and $\mathbf{F}^{2}_{L}$.
Comparison with vision-language models.   Our approach appears similar to recent vision-language models (VLMs) at the first glance, such as  Lin et al. (2023), FROMAGe (Koh et al., 2023), and LiMBeR (Merullo et al., 2023), where linear layers directly project visual features to the input space of LLMs. However, our approach is different, because the linear layer $\mathbf{F}_{L}^{1}$ does not necessarily align the visual representation $z$ into the language space. Concretely, this is reflected in three aspects: (1) Independence of visual pre-training. Our paradigm supports training-from-scratch without relying on pre-trained visual encoders like CLIP (Radford et al., 2021). (2) Independence of language. Our framework can function without language-based input or prompts, and it is applicable for general visual representation learning instead of only vision-language tasks. (3) Independence of transformer blocks. Previous VLMs treat an entire LLM as a coherent module, while our framework separates each transformer block as an independent layer for visual encoding.
Comparison with LLMs.   We substantially change the behaviors of LLM transformers, due to the distinct formats between visual and text data. (1) Attention mask. LLMs commonly utilize auto-regressive masks to mimic the order of text generation. However, the tokens in visual data come all at once, such as the image tokens of the cat (Fig. 1(a)). So we abandon auto-regressive attention masks and only use attention masks to indicate the padded tokens. (2) Positional embedding. The positional embedding in LLMs, e.g., rotary positional embedding (Su et al., 2021) in LLaMA, is not a common option for visual encoders. Therefore, we remove the positional embeddings of LLMs for simplicity and consistency with the original visual backbones. Considering the importance of attention masks and positional embeddings in transformers, it is surprising in hindsight that our framework has a positive influence on visual tasks.
@@ 4 Applicability of LLM Transformers for Visual Tasks
We instantiate our framework to various tasks and observe the wide applicability of pre-trained LLM transformers. Our experiments cover 2D (image classification) and 3D (point cloud classification), single-frame and multi-frame (action recognition), semantics and motion (motion forecasting), and tasks involving linguistic input or not (2D and 3D vision-language tasks). By default, we adopt the last transformer block from LLaMA-7B (Touvron et al., 2023). Our framework achieves consistent and significant improvements across these tasks following the standards of prior work. More analysis on our designs is in Sec. 5.
@@ 4.1 Image Classification
Image classification is the most common challenge for representation learning. We conduct experiments on ImageNet1k (Deng et al., 2009), and additionally evaluate on robustness benchmarks: corrupted images from ImageNet-C (Hendrycks & Dietterich, 2018), natural adversarial images from ImageNet-A (Hendrycks et al., 2021b), and out-of-distribution images from ImageNet-SK (Wang et al., 2019) and ImageNet-R (Hendrycks et al., 2021a).
Without loss of generality, we select ViT (Dosovitskiy et al., 2021) due to its wide use and native support for transformers. Following the notation in Eqn. 2, the encoder $\mathbf{F}_{E}$ is the set of self-attention transformer blocks and the decoder $\mathbf{F}_{D}$ denotes a linear classifier. An intuitive illustration is in Fig. 1(a). We train both the baseline ViT models and ViT+LLaMA from scratch following the same configuration of DeiT (Touvron et al., 2021). More details are in Sec. C.1.
The accuracy of ViT models consistently improves after incorporating the frozen LLaMA transformer block as in Table 1, including both the accuracy on clean ImageNet images and the robustness on corrupted or adversarial images. Our further experiments validate that the improvement is closely related to the LLM transformer instead of the sole consequence of an increased model capacity. Please refer to Sec. 5.1 and Sec. B.2 for details.
4.2 Point Cloud Classification
@@
Table 2: LLM transformer improves point cloud classification.
Point cloud classification handles a fundamentally different modality compared with images. The models predict labels by processing unordered 3D points and understanding the geometry. Our experiments cover two common datasets: ScanObjectNN (Uy et al., 2019) and ModelNet40 (Goyal et al., 2021). ScanObjectNN contains three splits: background (BG), foreground (OBJ), and clipped (T50) points. For ModelNet40, we experiment with different densities (1k, 4k, 8k) of points.
We adopt Point-BERT (Yu et al., 2021) and load its pre-trained parameters on ShapeNet (Chang et al., 2015). Then we append the LLaMA transformer after its final attention block before fine-tuning on point cloud classification datasets. Details are in Sec. C.2.
As shown in Table 2, our approach improves the accuracy for point cloud classification, further supporting the applicability of using a frozen LLM transformer as a visual encoding layer. Note that the accuracy slightly drops on ModelNet40 with 1k and 4k points, due to the saturation and $\sim$0.2variance of performance on ModelNet40 which is also analyzed in Ma et al. (2022). However, with an increased number of points (8k), the improvement of the LLaMA transformer is noticeable on ModelNet40. More importantly, on the more challenging ScanObjectNN, our approach improves the accuracy consistently and significantly. This experiment also shows that our framework is compatible with fine-tuning setups, in addition to training-from-scratch scenarios in Sec. 4.1.
@@ 4.3 Action Recognition
For the video modality, we apply the pre-trained LLM transformer block to action recognition, where the algorithm predicts the action labels of video clips. We choose the benchmark of “Something-something-v2” (SSv2) (Goyal et al., 2017b) for evaluation, because it highlights the challenge of understanding cross-frame movement, instead of relying on single-frame semantics.
@@
Table 3: Pre-trained LLM transformer improves action recognition on SSv2.
We follow VideoMAE (Tong et al., 2022) and adopt the simple yet effective ViT backbones. Different from patches of tokens in 2D images, the video tokens are cubes spanning both spatially and temporally. Identical to Fig. 1(a), we place the LLaMA transformer behind the last self-attention block in ViT. Our training setup also adopts the two-step practice in VideoMAE: (1) initialize ViT transformers from MAE (He et al., 2022) pre-training; (2) add the LLM transformer and then fine-tune on the SSv2 dataset using the same configuration of VideoMAE. More details are in Sec. C.3.
The LLaMA transformer enhances the accuracy for both ViT-S and ViT-B in Table 3, supporting the applicability of our framework for videos. To clarify, our baseline accuracy is lower than that reported in VideoMAE’s original paper, because VideoMAE used 32/64 GPUs to enable a larger batch size than our computational resources. Nonetheless, we control the settings identical between ViT and ViT-LLaMA for a fair comparison and indicate the positive effects of using LLM transformers.